OpenStudy (anonymous):

find a8 for the arithmatic sequence a1=6x-9, a2=5x+1

OpenStudy (precal):

is this one problem or 2 separate problems?

OpenStudy (anonymous):

one problem

OpenStudy (anonymous):

Find common difference: \(a_2 - a_1 = 5x + 1 - 6x - 9\) \(d = -x - 8\) Now \(a_8\) is given by: \[a_8 = a_1 + (8-1)(d) \implies a_8 = a_1 + 7d\] can you find \(a_8\) here ??

OpenStudy (anonymous):

maybe let me try

OpenStudy (anonymous):

Take your time..

OpenStudy (anonymous):

6x-9+7(-x-8) is that how I start it

OpenStudy (anonymous):

Distribute 7 to the brackets..

OpenStudy (anonymous):

Yes you are right Solve this and have faith on yourself.

OpenStudy (anonymous):

6x-9-7x-56

OpenStudy (anonymous):

Yes carry on..

OpenStudy (anonymous):

-x-65

OpenStudy (anonymous):

Well Done..

OpenStudy (anonymous):

is that the answer

OpenStudy (anonymous):

\(a_8 = -x - 65\) Yeah that looks good to me.

OpenStudy (anonymous):

the back of my book says the answer is 61 not 65

OpenStudy (anonymous):

We should then check it once again..

OpenStudy (anonymous):

Oh sorry I have mistaken above..

OpenStudy (anonymous):

with what

OpenStudy (anonymous):

\(a_2 - a_1 = 5x + 1 - 6x + 9\) \(d = -x + 10\) Now do the same.. Only value of d was wrong..

OpenStudy (anonymous):

why did you put plus 9

OpenStudy (anonymous):

\(a_8 = 6x - 9 + 7(-x + 10) \) \(a_8 = 6x - 9 - 7x + 70\)

OpenStudy (anonymous):

Do you know: -(-) becomes + ??

OpenStudy (anonymous):

its a plus 1 tho

OpenStudy (anonymous):

Where ??

OpenStudy (anonymous):

5x+1 and -6x-9

OpenStudy (anonymous):

+ implies +1 no doubt in this..

OpenStudy (anonymous):

Yes now subtract them..

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

\(5x + 9 - (6x - 9)\) = ??

OpenStudy (anonymous):

5x+1

OpenStudy (anonymous):

How??

OpenStudy (anonymous):

Oh sorry my mistake there..

OpenStudy (anonymous):

\(5x + 1 - (6x - 9) = ??\)

OpenStudy (anonymous):

-x-8

OpenStudy (anonymous):

No be careful..

OpenStudy (anonymous):

ohhh i see

OpenStudy (anonymous):

the negative before parenthasis

OpenStudy (anonymous):

Yes..

OpenStudy (anonymous):

Getting ??

OpenStudy (anonymous):

got it

OpenStudy (anonymous):

can u help with another

OpenStudy (anonymous):

Sure..

OpenStudy (anonymous):

find s12 a2=6, d=10

OpenStudy (anonymous):

what is s12 ??? 12th term ??

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

See, there is a general formula to find any term: \[\large a_n = a + (n-1)d\] Here: \[a_n = nth \; term\] a is first term of sequence, n is number of terms, and d is common difference..

OpenStudy (anonymous):

As here you want to find 12th term so here n = 12 a = 6 and d = 10 Can you find \(a_{12}\).. ??

OpenStudy (anonymous):

Oh sorry a2 is 6 and not a1..

OpenStudy (anonymous):

Firstly find a1 or a : \[a = a_2 - d\]

OpenStudy (anonymous):

@shanesullins are you getting or not?? Please do interact otherwise I am unable to help you..

OpenStudy (anonymous):

i got 6+(12-1)10 so far

OpenStudy (anonymous):

No here second term is 6 not first term.. You have to find first term here firstly.. First term you can find by: \[a = a_2 - d\] Can you find a here??

OpenStudy (anonymous):

-16

OpenStudy (anonymous):

Yes now use the formula..

OpenStudy (anonymous):

i got the wrong answer

OpenStudy (anonymous):

Let me try..

OpenStudy (anonymous):

\[a = a_2 - d \implies 6 - 10 = -4\]

OpenStudy (anonymous):

\[a_{12} = -4 + (12-1)(10) \implies -4 + 110 \implies \color{blue}{a_{12} = 106}\] Are not you getting this ??